OpenStudy (anonymous):

Find the sum of the series 6+2+2/62 +2/64 +2/66 +2/68 +....=282/35. I got a different answer using the formula S= a/(1-r), 6/(1-(1/36))= 216/35. What am I doing wrong?!

OpenStudy (anonymous):

It should be 6+2+2/6^2+2/6^4, etc

OpenStudy (anonymous):

forget about the 6 out front and add it later. this is a geometric series with \[a=2, r=\frac{1}{36}\] and you should get \[6+\frac{2}{1-\frac{1}{36}}=6+\frac{2}{\frac{35}{36}}=6+\frac{2\times 36}{35}\] whateverhtat is